OpenStudy (anonymous):

find the slope between points (2,2) and (3,2). A. unidentified slope B. 0 C. 1/2 D. - 1/2

jimthompson5910 (jim_thompson5910):

Hint: the y coordinates are the same, so you have a horizontal line

OpenStudy (anonymous):

so wouldjt it be C than?

jimthompson5910 (jim_thompson5910):

horizontal lines have a slope of 1/2 ??

OpenStudy (anonymous):

my bads its 0

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

so thats my answer than right?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

I have another question real quick: which point is contained in the graph of y>|x-2|-1?

jimthompson5910 (jim_thompson5910):

if the y coordinates are the same, then you have a horizontal line with a slope of 0

OpenStudy (anonymous):

I have another question real quick: which point is contained in the graph of y>|x-2|-1?

jimthompson5910 (jim_thompson5910):

ok what are your points

OpenStudy (anonymous):

well here are the multiple choices to chose from: A. 2,-1 B. 2, -2 C. 3,0 D. 2,0

OpenStudy (anonymous):

I got A for my answer

jimthompson5910 (jim_thompson5910):

let's find out if that's correct

jimthompson5910 (jim_thompson5910):

y>|x-2|-1 -1>|2-2|-1 ... plug in x = 2 and y = -1 -1 > |0| - 1 -1 > 0 - 1 -1 > -1 is that true?

OpenStudy (anonymous):

no lol maybe its B then

jimthompson5910 (jim_thompson5910):

plug in the coordinates and simplify

jimthompson5910 (jim_thompson5910):

if it's false, then it's not the answer

OpenStudy (anonymous):

hmm can you give me the problem to solve?

jimthompson5910 (jim_thompson5910):

y>|x-2|-1 -2>|2-2|-1 ... plug in x = 2 and y = -2 I'll let you finish

OpenStudy (anonymous):

i got no solution for that answer

jimthompson5910 (jim_thompson5910):

you should have gotten -2 > -1 which is also false

jimthompson5910 (jim_thompson5910):

so B isn't an answer either

OpenStudy (anonymous):

lol actually i think its C. kinda of a wild guess

jimthompson5910 (jim_thompson5910):

ok what do you get when you plug in x = 3 and y = 0

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

wait sorry hold on

OpenStudy (anonymous):

ya, I got no solution... 3>|0-2|-1

jimthompson5910 (jim_thompson5910):

what does the right side turn into?

OpenStudy (anonymous):

a possitive

OpenStudy (anonymous):

right?

jimthompson5910 (jim_thompson5910):

|0-2|-1 = |-2|-1 = ???

OpenStudy (anonymous):

always true

jimthompson5910 (jim_thompson5910):

|0-2|-1 = |-2|-1 = 2 - 1 = 1

OpenStudy (anonymous):

? so my answer would be D!!! right?

jimthompson5910 (jim_thompson5910):

oh wait, you plugged them in wrong but yes, your answer is D

jimthompson5910 (jim_thompson5910):

somehow you managed to eliminate C, but not sure how you did that

OpenStudy (anonymous):

i just solved wat u gave me:)

jimthompson5910 (jim_thompson5910):

i gotcha